# Additive Complementary Pairs of Codes

### v2

@@ 5 Constacyclic ${\rm I\!F}_{q}$-linear ACP of codes over ${\rm I\!F}_{q^{m}}$
This section focuses on counting formulas for constacyclic ${\rm I\!F}_{q}$-linear ACP of codes over ${\rm I\!F}_{q^{m}}$. Recall that two constacyclic ${\rm I\!F}_{q}$-linear codes $C$ and $D$ over ${\rm I\!F}_{q^{m}}$ form an ACP of codes if $C\bigoplus_{{\rm I\!F}_{q}}D={\rm I\!F}_{q^{m}}^{n}$. Let denote $q_{i}=q^{d_{i}}$. In the following theorem, we derive necessary and sufficient conditions under which a pair of constacyclic ${\rm I\!F}_{q}$-linear codes over ${\rm I\!F}_{q^{m}}$ form an ACP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1).
@@
Let $C$ and $D$ be two $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$, whose Canonical decompositions are given by (7) and (8), respectively. Then the pair $(C,D)$ forms an ACP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1) if and only if $C_{i}\bigoplus_{\mathcal{K}_{i}}D_{i}=\mathcal{I}_{i}$, for $0\leq i\leq r$ and $C_{r+k}\bigoplus_{\mathcal{K}_{r+k}}D_{r+k}=\mathcal{F}_{r_{+}k}$, for $1\leq k\leq s-r$.
@@
In order to study constacyclic ${\rm I\!F}_{q}$-linear ACP of codes over ${\rm I\!F}_{q^{m}}$, we first recall that, $|\mathcal{K}_{i}|=q^{d_{i}}$, where $|C^{q}_{1+ti}|=d_{i}$ for all $0\leq i\leq s$. Further $|\mathcal{I}_{i}|=\left(q^{m}\right)^{d_{i}}=|\mathcal{K}_{i}|^{m}$, for all $0\leq i\leq r$. Also, $|\mathcal{F}_{r+k}|=\left(q^{m}\right)^{d_{r+k}}=|\mathcal{K}_{r+k}|^{m}$, where $|C^{q}_{1+t(r+k)}|=d_{r+k}$ for all $1\leq r+k\leq s-r$. This implies that $\mathcal{I}_{i}$ is an $\mathcal{K}_{i}$-linear space of $\mathcal{K}_{i}^{m}$ for $0\leq i\leq r$ and $\mathcal{F}_{k+r}$ is an $\mathcal{K}_{k+r}$-linear space of $\mathcal{K}_{r+k}^{m}$ for $1\leq k\leq s-r$. It is well known that the number of ${\rm I\!F}_{q^{d_{i}}}$-subspace of ${\rm I\!F}_{q^{d_{i}}}^{m}$ is equal to ${\sum}\limits_{v=0}^{m}\left[\begin{array}[]{cc}m\\ v\end{array}\right]_{{q_{i}}}=\dfrac{(q_{i}^{m}-1)(q_{i}^{m}-q_{i})\cdots(q_{i}^{m}-q_{i}^{k-1})}{(q_{i}^{v}-1)(q_{i}^{v}-q_{i})\cdots(q_{i}^{v}-q_{i}^{v-1})}$. From the above discussion, we have the following theorem.
@@
By Theorem 5.1, the pair $(C,D)$ is an ACP if and only if $C_{i}\bigoplus_{\mathcal{K}_{i}}D_{i}=\mathcal{I}_{i}$, for $0\leq i\leq r$ and $C_{r+k}\bigoplus_{\mathcal{K}_{r+k}}D_{r+k}=\mathcal{F}_{r_{+}k}$ for $1\leq k\leq s-r$. Equivalently, $(C,D)$ is an ACP if and only if $(C_{i},D_{i})$ is an LCP for all $0\leq i\leq s$. If $(C_{i},D_{i})$ form an LCP for $0\leq i\leq s$, then the number of choices of $C_{i}$ is equal to
@@
and the number of choices of $D_{i}$ is equal to
@@
Therefore, the number of ACP of codes $(C,D)$ is
#$\#(C,D)={\prod}_{i=1}^{s}\left({\sum}\limits_{v=0}^{m}\left[\begin{array}[]{cc}m\\ v\end{array}\right]_{{q_{i}}}q_{i}^{v(m-v)}\right).$
@@
#$\#(C,D)={\prod}_{i=0}^{3}\left({\sum}\limits_{v=0}^{2}\left[\begin{array}[]{cc}2\\ v\end{array}\right]_{{3^{d_{i}}}}q_{i}^{v(2-v)}\right)={\prod}_{i=0}^{3}\left(2+2q_{i}\right)=4,40,896.$
6 An application on EAQEC codes
In this section, we construct entanglement assisted quantum error correcting (EAQEC) codes from ${\rm I\!F}_{q}$-linear ACP of codes. At first we present some basic concepts and notations on quantum codes. Let $\mathbb{C}^{q}$ denote the $q$-dimensional Hilbert space over the complex field $\mathbb{C}$ and $K$ be a finite field with characteristic $p$. Any $n$-qubit state is denoted by $|\mathbf{v}\rangle=\sum_{\mathbf{a}\in K^{n}}v_{\mathbf{a}}|\mathbf{a}\rangle$ where $v_{\mathbf{a}}\in\mathbb{C}$ with $\sum_{\mathbf{a}\in K^{n}}|v_{\mathbf{a}}|^{2}=1$ and the set $\{|\mathbf{a}\rangle=|a_{1}\rangle\otimes\cdots\otimes|a_{n}\rangle:(a_{1},\dots,a_{n})\in K^{n}\}$. For any two vectors $\mathbf{a}=(a_{1},\dots,a_{n})$ and $\mathbf{b}=(b_{1},\dots,b_{n})$ of $K^{n}$, the corresponding tensor products of $n$ error operators are given by $T(\mathbf{a})=T(a_{1})\otimes\cdots\otimes T(a_{n})$ and $R(\mathbf{b})=R(b_{1})\otimes\cdots\otimes R(b_{n})$. We define the error set $E_{n}:=\{\gamma^{i}T(\mathbf{a})R(\mathbf{b}):0\leq i\leq p-1;\mathbf{a},\mathbf{b}\in K^{n}\}$ where $\gamma$ is a complex primitive $p$-th root of unity. $E_{n}$ forms an error group.
Now, $T(\mathbf{a})|\mathbf{v}\rangle=|\mathbf{v}+\mathbf{a}\rangle$ and $R(\mathbf{b})|\mathbf{v}\rangle=\gamma^{tr_{\mathbb{F}_{q}/\mathbb{F}_{p}}\langle\mathbf{b},\mathbf{v}\rangle}|\mathbf{v}\rangle$. For any error $\mathbf{e}=\gamma^{i}T(\mathbf{a})R(\mathbf{b})$, we define the quantum weight of $\mathbf{e}$ as
$w_{Q}(\mathbf{e})=\{i:(a_{i},b_{i})\neq(0,0)\}.$
A positive integer $d$ is called the minimum distance of $q$-ary quantum code $Q$, if the following are satisfied
  • for any $|\mathbf{u}\rangle$, $|\mathbf{v}\rangle$ with $\langle\mathbf{u},\mathbf{v}\rangle=0$ such that $\langle\mathbf{u}|\mathbf{e}|\mathbf{v}\rangle=0$;
  • $\mathbf{e}\in E_{n}(d-1)$, where $E_{n}(i)=\{\mathbf{e}\in E_{n}:w_{Q}(\mathbf{e})\leq i\}$.
A $q$-ary quantum code with parameters length $n$, dimension $k$ and minimum distance $d$ is denoted by $[[n,k,d]]_{q}$. An EAQEC code represented as $[[n,k,d;c]]_{q}$, encodes $k$ logical qubits into $n$ physical qubits with $c$ copies of maximally entangled states. The performance of an EAQEC code is evaluated through its rate $\frac{k}{n}$ and net rate $\frac{k-c}{n}$, which can be positive, negative or zero. If the net rate of an EAQEC code is zero then it does not mean that no qubits are transmitted by this code. It actually means that a number of bits of entanglement is needed that is equal to the number of bits transmitted. The EAQEC codes with positive net rate are used in Quantum communication where as EAQECCs with negative net rate are used as Catalytic codes in Quantum computing (See [3, 4, 16] for more details). When $c=0$, the EAQEC code reduced to a standard stabilizer code. EAQEC codes can be regarded as generalized quantum codes.
A construction of EAQEC codes from the classical linear codes is provided by Wilde and Burne in [27].
Proposition 6.1.
[27, Corollary 1] Let $\mathcal{C}_{1}:[n,k_{1},d_{1}]$ and $\mathcal{C}_{2}:[n,k_{2},d_{2}]$ be two linear codes with parity check matrices $H_{1}$ and $H_{2}$, respectively. Then there exists an EAQEC code $[[n,k_{1}+k_{2}-n+c,min\{d_{1},d_{2}\};c]]_{q}$, where $c=rank(H_{1}H_{2}^{\top})$ is the required number of maximally entangled states.
Now, we will present a theorem which will be useful for designing EQECC codes later in this section. Towards this, for an ${\rm I\!F}_{q}$-linear additive code $C$ over ${\rm I\!F}_{q^{m}}$ of length $n$, we define
$Tr(C):=\{Tr(c):c\in C\}.$
Note that $Tr(C)$ is a linear code over ${\rm I\!F}_{q}$ of length $n$. We call $Tr(C)$ is a Trace code of $C$. Further, if $G$ is a generator matrix of $C$, then $Tr(G)$ (defined in Section 3)is a generator matrix of $Tr(C)$.
Theorem 6.1.
Let $C$ and $D$ be two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$. If the pair $(C,D)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then $(Tr(C),Tr(D))$ is an LCP of codes over ${\rm I\!F}_{q}$. In particular, if $C$ is an additive complementary dual (ACD) codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then $Tr(C)$ is an LCD codes over ${\rm I\!F}_{q}$.
Proof.
Let $G_{1}$ and $G_{2}$ be two generator matrices of $C$ and $D$, respectively. Since, the pair $(C,D)$ is an ACP of codes over ${\rm I\!F}_{q^{m}}$ (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1), then by Theorem 3.3, we obtain that $rank\left(Tr\left(\begin{array}[]{cc}G_{1}\\ G_{2}\end{array}\right)\right)=n.$ This gives that $rank\left(\begin{array}[]{cc}Tr(G_{1})\\ Tr(G_{2})\end{array}\right)=n$. Since $Tr(G_{1})$ and $Tr(G_{2})$ are generator matrices of $Tr(C)$ and $Tr(D)$, respectively, from Proposition 4.1, we get $(Tr(C),Tr(D))$ is an LCP of codes over ${\rm I\!F}_{q}$. ∎
We have the following theorem deducted from Proposition 6.1.
Theorem 6.2.
Let $C$ and $D$ be two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$ such that $Tr(C):=[n,k_{1},d_{1}]$ and $Tr(D):=[n,k_{2},d_{2}]$. Let parity check matrices of $C$ and $D$ be $H_{1}$ and $H_{2}$, respectively . Then there exists an EAQEC code $[[n,k_{1}+k_{2}-n+c,min\{d_{1},d_{2}\};c]]_{q}$ where $c=rank(Tr(H_{1})Tr(H_{2})^{\top})$ is the required number of maximally entangled states.
Corollary 6.3.
Let $(C,D)$ be an ACP of codes of ${\rm I\!F}_{q^{m}}^{n}$ such that $Tr(C):=[n,k_{1},d_{1}]$ and $Tr(D):=[n,k_{2},d_{2}]$. Let parity check matrices of $C$ and $D$ be $H_{1}$ and $H_{2}$, respectively. Then there exists an EAQEC code $[[n,c,min\{d_{1},d_{2}\};c]]_{q}$ where $c=rank(Tr(H_{1})Tr(H_{2})^{\top})$ is the required number of maximally entangled states. Moreover, if $D=C^{\perp}$ with parity check matrix $H$, then there exists an EAQEC code $[[n,k_{1},d_{1};n-k_{1}]]_{q}$.
It is well known that $rank(Tr(H)Tr(H)^{\top})=\dim(Tr(C^{\perp}))-dim(hull(Tr(C)))$ where $H$ is a parity check matrix of $C$.
Corollary 6.4.
Let $C$ be a ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$ such that $Tr(C):=[n,k,d]$. Let $H$ be a parity check matrix of $C$. Then there exists an EAQEC code $[[n,k-\dim(hull(Tr(C))),d_{H};c]]_{q}$ where $c=rank(Tr(H)Tr(H)^{\top})$ is the required number of maximally entangled states.
@@ 6 An application on EAQEC codes
In Table 1, we present some examples of LCD codes corresponding additive code from Theorem 6.1. In Table 1, we we present some examples of EAQEC codes from Trace code of additive code from Corollary 6.4.
@@
No. Generator of additive codes Trace Codes EAQEC codes from Existing
Corollary 6.4
$[[n,k-dim(hull(\mathcal{C})),d_{H};$ EAQEC codes
$n-k-dim(hull(\mathcal{C}))]]_{2}$
1 $\left(\begin{array}[]{cccccc}\omega&\omega^{2}&0\\ 0&\omega&\omega^{2}\end{array}\right)$ $[3,~{}2,~{}2]$ $[[3,~{}2,~{}2;~{}1]]_{2}$ $[[3,~{}2,~{}2;~{}1]]_{2}$[15]
2 $\left(\begin{array}[]{ccccccc}\omega&1&0&1&\omega&\omega^{2}\\ 0&\omega&0&1&1&\omega\\ 1&1&\omega^{2}&1&\omega&1\\ 0&1&0&\omega^{2}&\omega&\omega^{2}\\ \end{array}\right)$ $[6,~{}4,~{}2]$ $[[6,~{}2,~{}2;~{}0]]_{2}$ $[[6,~{}2,~{}2;~{}0]]_{2}$[15]
3 $\left(\begin{array}[]{ccccccc}\omega^{2}&1&0&1&1&\omega^{2}\\ 0&\omega&0&1&1&\omega\\ 1&1&\omega^{2}&1&1&\omega^{2}\\ 0&1&0&\omega^{2}&0&\omega\\ 0&0&0&0&\omega&\omega^{2}\\ \end{array}\right)$ $[6,~{}5,~{}2]$ $[[6,~{}4,~{}2;~{}0]]_{2}$ $[[6,~{}2,~{}2;~{}0]]_{2}$[15]
4 $\left(\begin{array}[]{cccccccc}\omega&1&0&1&1&\omega^{2}&\omega\\ 0&\omega&0&0&\omega&1&\omega\\ 1&1&\omega^{2}&1&\omega^{2}&\omega&1\\ 0&1&0&\omega^{2}&\omega&\omega&\omega^{2}\\ \end{array}\right)$ $[7,~{}4,~{}3]$ $[[7,~{}1,~{}3;~{}0]]_{2}$ $[[7,~{}1,~{}3;~{}0]]_{2}$[15]
6 $\left(\begin{array}[]{ccccccccc}\omega&\omega&0&\omega&\omega&0&\omega&\omega&0\\ 0&\omega&\omega&0&\omega&\omega&0&\omega&\omega\end{array}\right)$ $[9,~{}2,~{}6]$ $[[9,~{}2,~{}6;~{}7]]_{2}$ $[[9,~{}2,~{}6;~{}7]]_{2}$[19]
7 $\left(\begin{array}[]{ccccccccccccccc}\omega&\omega^{2}&0&\omega&\omega^{2}&0&\omega&\omega^{2}&0&\omega&\omega^{2}&0&\omega&\omega^{2}&0\\ 0&\omega&\omega^{2}&0&\omega&\omega^{2}&0&\omega&\omega^{2}&0&\omega&\omega^{2}&0&\omega&\omega^{2}\\ \end{array}\right)$ $[15,~{}2,~{}10]$ $[[15,~{}2,~{}10;~{}13]]_{2}$ $[[15,~{}2,~{}10;~{}13]]_{2}$[19]
Table 2: Some EAQEC codes
@@ 7 Conclusion
In this work, we study the characterization and construction of an ACP of codes of length $n$ over ${\rm I\!F}_{q^{m}}$. Moreover, we obtain a condition for an additive pairs of codes to be an ACP of codes. Furthermore, we provide a necessary and sufficient condition for an ACP of codes. Finally, we obtain an ACP of codes from given linear complementary pairs of codes. In addition, as an immediate consequence, we exhibited some optimal binary LCD codes and corresponding EAQCC codes.
Acknowledgement
The authors would like to thank Markus Grassl for their comments and suggestions which us to improve our manuscript.

### v3

@@ 5 Constacyclic ${\rm I\!F}_{q}$-linear ACP of codes over ${\rm I\!F}_{q^{m}}$
This section focuses on counting formulas for constacyclic ${\rm I\!F}_{q}$-linear ACP of codes over ${\rm I\!F}_{q^{m}}$. Recall that two constacyclic ${\rm I\!F}_{q}$-linear codes $\mathcal{C}$ and $\mathcal{D}$ over ${\rm I\!F}_{q^{m}}$ form an ACP of codes if $\mathcal{C}\bigoplus_{{\rm I\!F}_{q}}\mathcal{D}={\rm I\!F}_{q^{m}}^{n}$. Let denote $q_{i}=q^{d_{i}}$. In the following theorem, we derive necessary and sufficient conditions under which a pair of constacyclic ${\rm I\!F}_{q}$-linear codes over ${\rm I\!F}_{q^{m}}$ form an ACP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1).
@@
Let $\mathcal{C}$ and $\mathcal{D}$ be two $\lambda$-constacyclic ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$, whose Canonical decompositions are given by (5) and (6), respectively. Then the pair $(\mathcal{C},\mathcal{D})$ forms an ACP of codes (using identity $\pi$, $\sigma$ and $M$ in Proposition 4.1) if and only if $\mathcal{C}_{i}\bigoplus_{\mathcal{K}_{i}}\mathcal{D}_{i}=\mathcal{I}_{i}$, for $0\leq i\leq r$ and $\mathcal{C}_{r+k}\bigoplus_{\mathcal{K}_{r+k}}\mathcal{D}_{r+k}=\mathcal{F}_{r_{+}k}$, for $1\leq k\leq s-r$.
@@
In order to study constacyclic ${\rm I\!F}_{q}$-linear ACP of codes over ${\rm I\!F}_{q^{m}}$, we first recall that, $|\mathcal{K}_{i}|=q^{d_{i}}$, where $|\mathcal{C}^{q}_{1+ti}|=d_{i}$ for all $0\leq i\leq s$. Further $|\mathcal{I}_{i}|=\left(q^{m}\right)^{d_{i}}=|\mathcal{K}_{i}|^{m}$, for all $0\leq i\leq r$. Also, $|\mathcal{F}_{r+k}|=\left(q^{m}\right)^{d_{r+k}}=|\mathcal{K}_{r+k}|^{m}$, where $|\mathcal{C}^{q}_{1+t(r+k)}|=d_{r+k}$ for all $1\leq r+k\leq s-r$. This implies that $\mathcal{I}_{i}$ is an $\mathcal{K}_{i}$-linear space of $\mathcal{K}_{i}^{m}$ for $0\leq i\leq r$ and $\mathcal{F}_{k+r}$ is an $\mathcal{K}_{k+r}$-linear space of $\mathcal{K}_{r+k}^{m}$ for $1\leq k\leq s-r$. It is well known that the number of ${\rm I\!F}_{q^{d_{i}}}$-subspace of ${\rm I\!F}_{q^{d_{i}}}^{m}$ is equal to ${\sum}\limits_{v=0}^{m}\left[\begin{array}[]{cc}m\\ v\end{array}\right]_{{q_{i}}}=\dfrac{(q_{i}^{m}-1)(q_{i}^{m}-q_{i})\cdots(q_{i}^{m}-q_{i}^{k-1})}{(q_{i}^{v}-1)(q_{i}^{v}-q_{i})\cdots(q_{i}^{v}-q_{i}^{v-1})}$. From the above discussion, we have the following theorem.
@@
By Theorem 5.1, the pair $(\mathcal{C},\mathcal{D})$ is an ACP if and only if $\mathcal{C}_{i}\bigoplus_{\mathcal{K}_{i}}\mathcal{D}_{i}=\mathcal{I}_{i}$, for $0\leq i\leq r$ and $\mathcal{C}_{r+k}\bigoplus_{\mathcal{K}_{r+k}}\mathcal{D}_{r+k}=\mathcal{F}_{r_{+}k}$ for $1\leq k\leq s-r$. Equivalently, $(\mathcal{C},\mathcal{D})$ is an ACP if and only if $(\mathcal{C}_{i},\mathcal{D}_{i})$ is an LCP for all $0\leq i\leq s$. If $(\mathcal{C}_{i},\mathcal{D}_{i})$ form an LCP for $0\leq i\leq s$, then the number of choices of $\mathcal{C}_{i}$ is equal to
@@
and the number of choices of $\mathcal{D}_{i}$ is equal to
@@
Therefore, the number of ACP of codes $(\mathcal{C},\mathcal{D})$ is
#$\#(\mathcal{C},\mathcal{D})={\prod}_{i=1}^{s}\left({\sum}\limits_{v=0}^{m}\left[\begin{array}[]{cc}m\\ v\end{array}\right]_{{q_{i}}}q_{i}^{v(m-v)}\right).$
@@
#$\#(\mathcal{C},\mathcal{D})={\prod}_{i=0}^{3}\left({\sum}\limits_{v=0}^{2}\left[\begin{array}[]{cc}2\\ v\end{array}\right]_{{3^{d_{i}}}}q_{i}^{v(2-v)}\right)={\prod}_{i=0}^{3}\left(2+2q_{i}\right)=4,40,896.$
@@
In Table 1, we present some examples of LCD codes corresponding to additive code from Theorem 4.1.
@@ 6 Conclusion
The main contribution of this paper is that a very general type of ACP codes is characterized and constructed. Moreover, we obtained a condition for an additive pair of codes to be an ACP of codes. Furthermore, we provide a necessary and sufficient condition for an ACP of codes. Finally, we obtained an ACP of codes from given linear complementary pairs of codes.